Let x and y equal real numbers; (x+3yi)-(9-25i)=-5+4i
solve for x and y

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Since



We can write:





John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Let x and y equal real numbers;
solve for x and y
(x+3yi)-(9-25i)=-5+4i
---
(x-9)+(3y+25)i = -5+4i
----
Equate the Real Parts:
Equate the Imaginary Parts:
x-9 = -5
3y+25 = 4
---------------------
Solve for "x" and for "y":
x = 4
y = -7
=============
Cheers,
Stan H.